OpenStudy (anonymous):

Which equation represents the line passing through the point (4,-16) and is perpendicular to the line y=-2/3 x+8?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (abmon98):

-2/3*m=-1 m=3/2 y+16=3/2(x-3) 2y+32=3x-9 y=3/2x-20.5

OpenStudy (anonymous):

@skullpatroll i don't want you here.

OpenStudy (abmon98):

OpiGeode is the answer right?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

my gf said it was 2/3 i'm not sure.

OpenStudy (anonymous):

did she have the answer's because Abmon's work looks good to me